OpenStudy (anonymous):

Julie is a math tutor in the academic center. She charges $210 for the first 7 hours of tutoring and $20 for each additional hour of tutoring. If she earned $410 tutoring 1 student, how many total hours did she spent tutoring the student? Which of these equations represents this problem? A) h = 7x ( 410 + 210)/20 B) H=7-(410+210)/20 C) H=7-(410-210)/20 D)H=7+(410-210)/20

OpenStudy (anonymous):

h = 7 + (410-210)/20

OpenStudy (ybarrap):

She made $210 for the 1st 7 hours and 20 dollars for each additional hours, so, if h represents the number of additional hours beyond 7 hours she tutors, we have, $$ 410=210 + 20h\\ h=\cfrac{410-210}{20}\\ $$ So the total number of hours would be 7 + h.